USEU How many I atoms are there in 5.45 moles of I? atoms Submit Answer Retry Entire Group 6 mo
Answer #1

Answer:

Avogadro's Hypothesis: 1 mole of the substance contains 6.022*1023 particles of the substance.

Hence, we write,

1 mole of I = 6.022*1023 of I atoms.

\Rightarrow 5.45 mol of I = 5.45 * 6.022*1023 of I atoms = 3.282*1024 of I atoms.

5.45 mol of I contains 3.282*1024 of I atoms.

====================XXXXXXXX==============